a pharmacist is to prepare 15 milliliters of special eye drops. The eye drop solution must have 2% active ingredient, but the pharmacist only has 10% solution and 1% solution in stock. How much of each type of solution should be used to fill the prescription?
Answer by mananth(16949)

----------------- percent ----------------quantity
Eye solution type I 10 ----------------x ml
Eye solution type II 1 ----------------15-x ml
Mixture 2.00% ---------------- 15

10x+(15-x)=2*15

10x+15-x=30
10x-x=30-15
9x=15
/9
x=1.67 ml 10.00% Eye solution type I
13.33 ml 1.00% Eye solution type II
m.ananth@hotmail.ca